# Threshold Graphs are Globally Synchronizing

### v4

@@ 6.1. A Prototype Example
We next verify that the sets $Q,S_{1},S_{2}$, and $P$ satisfy the assumptions of Lemma 5.8. First, for every node $i$ in $Q$, its neighborhood is contained in $S_{1}$; indeed, these nodes are isolated at this stage and connect only to the dominating nodes that are added later. Second, the sets $S_{1}$ and $S_{2}$ together induce a clique. Third, for any node in $S_{1}\cup S_{2}$, its neighbors consist exactly of all other nodes in $S_{1}\cup S_{2}$, all nodes in $P$, and possibly some nodes in $Q$. Hence all conditions of Lemma 5.8 are satisfied, and we conclude that $S_{1}$ and $S_{2}$ must synchronize. In particular, the block $U_{k-m}$ synchronizes with $U_{k-m+1}\uplus U_{k-m+2}\uplus\cdots\uplus U_{k}$.
By induction, all nodes synchronize at $\boldsymbol{\theta}$. Therefore, every second-order stationary point is a synchronous state. It follows that every connected threshold graph is globally synchronizing. ∎
7. Discussion: Beyond Density-Based Guarantees for Global Synchronization
As noted in the introduction, previous guarantees for global synchronization have largely focused on minimum-degree conditions: that is, requiring every node to be connected to more than a fixed fraction ($\geq 75\%$) of the remaining vertices. In this regime, graphs are necessarily highly dense, and their degree sequences are relatively homogeneous. This naturally raises the question:
Does a graph that exhibits global synchronization necessarily have high edge density or a homogeneous degree sequence (i.e., with relatively small variation in node degrees)?
In this section, we provide a negative answer to this question [^5] Of course, trees provide trivial examples. by exploiting two structural properties of threshold graphs, which were shown in the previous sections to be globally synchronizing. More precisely, we show that threshold graphs realize every admissible edge density among connected graphs. Moreover, their degree sequences are highly heterogeneous: among all connected graphs with a fixed edge density, threshold graphs attain the maximally left-skewed degree sequence when ordered in nonincreasing order.
@@ 7.1. Threshold Graphs Realize All Densities
Note that the edge density of a graph $G$ on $n$ vertices takes values in
@@
and for connected graphs
$\mathrm{density}(G)\in\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{n-1,\ldots,\binom{n}{2}\}\,\right\}$. (21)
If the graph is not connected, to realize any density in $\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{0,\ldots,n-2\}\,\right\}$, construct a threshold graph on $n$ vertices whose n1n-1-length sequence contains $1$ at the $m$-th position.
001
Figure 15. Illustration of the threshold graph construction. A single 11 at the third position of the generating sequence adds a dominating vertex, which contributes three incident edges shown as curved arcs.
The following lemma shows that every admissible edge density in the range (21) is realized by a connected threshold graph on $n$ vertices.
@@
Fix $n\geq 2$. For any integer $m\in[\,n-1,\binom{n}{2}\,$], there exists a connected threshold graph on $n$ vertices with exactly $m$ edges.
@@
Since the graph is connected, the last vertex added in the construction of a threshold graph must be a dominating vertex, contributing n1n-1 edges. To obtain a graph with exactly $m$ edges, the vertices added before the last one must therefore contribute a total of m(n1)m-(n-1) edges. Since $m\in\{n-1,\ldots,\binom{n}{2}\}$, it follows that
$m-(n-1)\in\left\{0,1,\ldots,\binom{n}{2}-(n-1)\right\}=\left\{0,1,\ldots,\frac{(n-2)(n-1)}{2}\right\}.$
Edges are created only when dominating vertices are added during the construction. We fix the following convention: starting from an initial vertex, we index the construction sequence by counting only the dominating vertices added thereafter. Under this convention, a dominating vertex added at position $p$ contributes exactly $p$ new edges, as it connects to all previously existing vertices. Hence the possible edge contributions of dominating vertices, except for the last one added in the construction sequence, take values in $\{1,2,\ldots,n-2\}$.
Therefore, the problem reduces to determining at which positions the dominating vertices except for the last one are added so that they collectively contribute m(n1)m-(n-1) edges. More precisely, it suffices to prove the following combinatorial statement:
For any integer
$k\in\left\{0,1,\ldots,\frac{(n-2)(n-1)}{2}\right\},$
there exists a subset
$A\subseteq\{1,2,\ldots,n-2\}\quad\textit{such that}\quad\sum_{a\in A}a=k.$
We prove this claim by induction on $n$.
Base case. For $n=3$, we have $\{1,2,\ldots,n-2\}=\{1\}$ and $k\in\{0,1\}$. If $k=0$, take $A=\emptyset$; if $k=1$, take $A=\{1\}$.
Induction hypothesis. Assume that for any
$k\in\left\{0,1,2,\ldots,\frac{(n-3)(n-2)}{2}\right\},$
there exists a subset
$A^{\prime}_{k}\subseteq\{1,2,\ldots,n-3\}\quad\text{such that}\quad\sum_{a\in A^{\prime}_{k}}a=k.$
Induction step. We now consider any
$k\in\left\{0,1,2,\ldots,\frac{(n-2)(n-1)}{2}\right\}.$
Case 1. If
$k\leq\frac{(n-3)(n-2)}{2},$
then by the induction hypothesis there exists $A^{\prime}_{k}\subseteq\{1,2,\ldots,n-3\}$ such that $\sum_{a\in A^{\prime}_{k}}a=k$. Setting $A:=A^{\prime}_{k}$ yields the desired subset.
Case 2. If
$k>\frac{(n-3)(n-2)}{2},$
then using only values in $\{1,2,\ldots,n-3\}$ cannot achieve the sum $k$, since the maximal possible sum is
$\sum_{i=1}^{n-3}i=\frac{(n-2)(n-3)}{2},$
which is strictly smaller than $k$. Hence the element n2n-2 must be included in $A$.
It then remains to find a subset
$A^{\prime\prime}\subseteq\{1,2,\ldots,n-3\}\quad\text{such that}\quad\sum_{a\in A^{\prime\prime}}a=k-(n-2).$
Since
$k-(n-2)\in\left\{\frac{(n-2)(n-5)}{2}+1,\ldots,\frac{(n-2)(n-3)}{2}\right\}\subseteq\left\{0,1,\ldots,\frac{(n-3)(n-2)}{2}\right\},$
the induction hypothesis guarantees the existence of such a subset; in particular, we may take
$A^{\prime\prime}=A^{\prime}_{\,k-(n-2)}.$
Setting $A:=A^{\prime\prime}\cup\{n-2\}$ then yields $\sum_{a\in A}a=k$.
This completes the induction and hence the proof. ∎
@@ 7.2. Degree Sequence of Threshold Graphs
One way to compare the “shape” of the degree sequences of two graphs with the same edge density is through majorization, which provides a partial order quantifying how spread out or concentrated the degrees are. More specifically, let $G$ and $G^{\prime}$ be two graphs with the same edge density, and let
@@
As shown in the previous section, for any admissible edge density there exists at least one (not necessarily unique) threshold graph realizing that density. Moreover, Theorem 7.2 tells that, among all graphs realizing a given edge density, threshold graphs are exactly those whose degree sequences are maximal with respect to majorization. In this sense, threshold graphs are always the most left-skewed among graphs with a fixed edge density, since their degree sequences are maximal under majorization, meaning that the degree mass is as concentrated as possible on high-degree vertices.
@@ 8. Summary and Open Problems
In this work, we present a class of globally synchronizing graphs — threshold graphs — whose density vary from $2/n$ to $1$. Our proof crucially exploits the constructive nature of threshold graphs: they form the smallest graph class closed under the addition of isolated nodes and universal nodes. This closure property is essential to our inductive argument, as it allows us to show that global synchronization is preserved when starting from a threshold graph and adding either an isolated node or a universal node.

### v5

@@ 6.1. A Prototype Example
We now verify that the assumptions of Lemma 5.9 are satisfied by the above decomposition, which follows directly from the threshold construction.
By construction, the vertices in $Q$ correspond to the most recent 0-blocks preceding $S_{1}$. Hence, every vertex in $Q$ is adjacent only to vertices in $S_{1}$. Moreover, since $S_{1}$ and $S_{2}$ consist of consecutive dominating blocks, the induced subgraph on $S_{1}\cup S_{2}$ is a clique. Finally, for each $i\in S_{1}$, its neighborhood can be decomposed as
$N(i)=P\;\uplus\;\bigl((S_{1}\cup S_{2})\setminus\{i\}\bigr)\;\uplus\;N_{Q}(i),$
where $P$ denotes the set of vertices added earlier in the construction. Hence, all the assumptions of Lemma 5.9 are satisfied, and we conclude that the sets $S_{1}$ and $S_{2}$ synchronize. In other words, the last $m+1$ blocks of $1$’s $U_{k-m},U_{k-m+1},\cdots,U_{k}$ synchronize.
Therefore, every second-order stationary point is a synchronous state, and it follows that $G$ is globally synchronizing. ∎
7. Discussion: Beyond Density-Based Guarantees
As noted in the introduction, previous guarantees for global synchronization have largely focused on minimum-degree conditions: that is, requiring every node to be connected to more than a fixed fraction of the remaining vertices. In this regime, graphs are necessarily highly dense, and their degree sequences are relatively homogeneous. This naturally raises the question:
Does a graph that exhibits global synchronization necessarily have high edge density or a nearly regular degree sequence (i.e., with small degree variability)?
In this section, we provide a negative answer to this question [^4] Of course, trees provide trivial examples. by showing that threshold graphs realize every admissible edge density, and that their degree sequences are extremal in the sense that they majorize any other degree sequence.
@@ 7.1. Threshold Graphs Realize All Densities
The edge density of a graph $G$ on $n$ vertices takes values in
@@
Moreover, for connected graphs, the edge density takes values corresponding to
$m\in\{n-1,\ldots,\binom{n}{2}\}$.
The following lemma shows that every admissible edge density can be realized by a threshold graph.
@@
For any pair of integers $n>0$ and $0\leq m\leq\binom{n}{2}$, there is a threshold graph with $n$ vertices and $m$ edges. If $m\geq n-1$, there is a connected threshold graph with $n$ vertices and $m$ edges.
@@
The proof is by induction over $n$. For the base case, note that the only graph with $n=1$ vertex has $0=m=\binom{n}{2}$ edges and is connected by definition. For $n>1$, we consider two cases:
  • If m<n1m<n-1, the induction hypothesis implies that there is a threshold graph with $n-1\geq 1$ vertices and $m<n-1\leq\binom{n-1}{2}$ edges to which an isolated vertex can be added to obtain a threshold graph with $n$ vertices and $m$ edges.
  • Otherwise, the induction hypothesis implies that there is a threshold graph with $m-(n-1)\leq\binom{n-1}{2}$ edges to which a universal vertex can be added to obtain a connected threshold graph with $n$ vertices and $m$ edges.
@@ 7.2. Degree Sequence of Threshold Graphs
One way to compare how spread out the degree sequences of two graphs are is through majorization. Let $G$ and $G^{\prime}$ be two graphs with the same edge density, and let
@@
This theorem formalizes the intuition that threshold graphs are extremal, in the sense that their degree mass is concentrated as much as possible on high-degree vertices.
@@ 8. Summary and Open Problems
In this work, we identify threshold graphs as a class of interaction networks that are globally synchronizing, despite spanning the full range of admissible edge densities from $2/n$ to $1$. This significantly extends existing sufficient conditions for global synchronization, which are largely restricted to graphs with high minimum degree or uniformly dense connectivity.
Our analysis crucially exploits the constructive nature of threshold graphs: they form the smallest graph class closed under the successive addition of isolated and universal nodes. This closure property enables an inductive approach, whereby global synchronization is shown to be preserved when starting from a smaller synchronizing threshold graph and adding either an isolated node or a universal node. From a geometric perspective, this reveals a local-to-global propagation mechanism, whereby synchronization constraints imposed on a small subset of nodes progressively force synchrony across the entire network.